OpenStudy (anonymous):

How would I go about factoring something like x^3+x-2?

OpenStudy (anonymous):

\[x^3+x-2\]

OpenStudy (loser66):

\[= (x+1)^2(x-2)\]

OpenStudy (anonymous):

how do you get there?

OpenStudy (loser66):

+x^2 -x^2 then, factor

OpenStudy (loser66):

got it?

OpenStudy (anonymous):

if i distribute that out i get x^3-3x-2

OpenStudy (loser66):

show me yours step

OpenStudy (anonymous):

\[(x+1)^2= x^2+2x+1\] \[(x^2+2x+1)(x-2)=x^3+2x^2+x-2x^2-4x-2\] \[=x^3-3x-2\]

OpenStudy (loser66):

so, go backward you have mine. that what you need: factor!!

OpenStudy (loser66):

now, I rearrange it for you

OpenStudy (loser66):

from yours: \[ x^3 +x -2 = x^3 +x^2-x^2+x-2\]

OpenStudy (loser66):

\[(x^3 +x^2) - (x^2-x+2)= x^2(x+1)-(x-2)(x+1)\]

OpenStudy (anonymous):

oh i see it

OpenStudy (loser66):

\[=(x+1)(x^2-(x-2))=(x+1)(x^2-x+2)\]

OpenStudy (anonymous):

thank you

OpenStudy (loser66):

good yw